# Towards automated formal security analysis of SAML V2.0 Web Browser SSO standard — the POST/Artifact use case

### v1

@@ III-A Methodology and Model Variants
¡AuthnRequest¿  ¡Subject¿
Specifies the subject of the authentication request or assertion, which identifies the principal that is the subject of the request or assertion.
¡AuthnRequest¿  ¡Version¿
Specifies the version of the SAML protocol.
¡AuthnRequest¿  ¡IssueInstant¿
Specifies time at which the request or assertion was issued.
¡AuthnRequest¿  ¡ForceAuthn¿
Indicates whether the identity provider must authenticate the principal even if the principal has an existing session with the identity provider.
¡AuthnRequest¿  ¡IsPassive¿
Indicates whether the identity provider must not interact with the principal, such as by prompting the principal for credentials, during the authentication process.
¡Response¿  ¡EncryptedAssertion¿
Represents an assertion that has been encrypted to protect its confidentiality.
¡Response¿  ¡Assertion¿   ¡AuthnStatement¿
Contains information about the authentication of a principal, including the time of authentication and the method of authentication used.
¡Response¿  ¡Assertion¿   ¡Subject¿    ¡SubjectConfirmationData¿     ¡OneTimeUse¿
Indicates whether the authentication request or assertion can only be used once, to prevent replay attacks.
¡Response¿  ¡Assertion¿   ¡Subject¿    ¡SubjectConfirmationData¿     ¡NotOnOrAfter¿
Specifies the time after which the authentication request or assertion is no longer valid.
@@
TABLE I: Elements that are not implemented in our model which are required by SAML standard or/and relevant to the security properties of protocol.
The other source of difficulty when converting the standard to the formal model is the lack of specification of various mechanism. Here, we tried to asses if the under-specified mechanism plays a purely functional role or if it could potentially have an impact on the security properties of the overall system. In the latter case, we include the mechanism in the list of sensitive under-specified mechanisms. For example, the SAML standard specifies that “dereferencing of the artifact using the Artifact Resolution profile MUST be mutually authenticated, integrity protected, and confidential“ [25], but does not explicitly specify the mechanisms to achieve these properties.
When deciding how exactly to model an under-specified mechanism, we considered a straightforward canonical implementation and, additionally examined various SAML implementations to identify how are those under-specified mechanisms implemented in practice.
After identifying the sensitive optional features and sensitive under-specified mechanisms, we decided to pick three of these and model each in two different ways yielding eight different protocol variants. Since our goal was to investigate if and how small changes impact the security of the complete system, we picked the features and implementations that seemed most likely to have a significant impact, but are also illustrative of kinds choices and potential pitfalls in building a standards-compliant SAML deployment.
@@
SAML specification mandates the use identifiers for many SAML messages. The specification requires that “any party that assigns an identifier MUST ensure that there is negligible probability any other party will accidentally assign the same identifier to a different data object” [21]. Notice that this does not imply that identifiers must be generated via a cryptographic pseudorandom number generator (PRNG) — only that the probability of collisions between identifiers generated by honest protocol parties is negligibly low. Clearly, this is a sensitive under-specified mechanisms since SAML identifiers are used to match responses to the original requests.
@@
We model two different behaviours corresponding to cryptographically strong and weak random identifiers:
@@ III-B Communication channels
Our model of TLS communication — given in Listing 1 — is similar to one in [35], except that we use ordinary network channels for communication and wrap TLS messages inside private function names. Using ordinary network channels significantly reduced precomputation time while wrapping TLS messages inside private functions makes it impossible for adversary to read the messages. Furthermore, we do not require channel IDs as in [35] since TLS sessions in SAML SSO are short-lived (only one request and response).
1functions: https_request/3 [private],
2           https_response/2 [private]
3rule Client_Starts_TLS:
4[
5  Fr(~tls_session_id),
6  !Is_Agent(\$Client),
7  !Is_Agent(\$Server)
8]
9--[
10  Client_Starts_TLS(\$Client, \$Server, ~tls_session_id)
11]->
12[
13  New_Client_TLS_Session(
14      \$Client,\$Server, ~tls_session_id),
15  New_Server_TLS_Session(\$Server, ~tls_session_id)
16]
17
18rule Sample_TLS_Request:
19[
20  New_Client_TLS_Session(
21      \$Client, \$Server, ~tls_session_id)
22]
23--[]->
24[
25  Out(https_request(\$Server, ~tls_session_id, request))
26]
27
28rule Sample_TLS_Response:
29[
30  New_Server_TLS_Session(\$Server, ~tls_session_id),
31  In(https_request(\$Server, ~tls_session_id, request))
32]
33--[]->
34[
35  Out(https_response(~tls_session_id, response))
36]
Listing 1: TLS Channel rules
TLS session is established between two agents (the notion of agents is precisely described in the following section III-C). For every TLS request towards agent acting as server S, a new session is created and labeled with unique tls_session_id. Since we model one-way authenticated TLS, any agent can act as client and initiate such session. However, only agent S can access the data in the request, and create a response bound to the same tls_session_id Such response can be accessed only by the agent that initiated the TLS session. Note that the agents name S doubles as its TLS identity, in practice this can be, for example, its common name or a pinned certificate.
III-C System setup
[…]
[t](botskip=0pt, midskip=0pt)[scale=0.75]fig/hartl2.pdf Prerequisites Rules state machine. Initialized agent can become one of three honest parties or become compromised. Clients (and the adversary) can register resources at SPs (honest or compromised). Client (and the adversary) can register at IdPs (honest or compromised). Finally, any two SAML parties (SP and IdP) can establish trust. In our model, the system consists of a number of agents. There are three kinds of honest agents – the SPs and IdPs whose behaviour is defined by the SAML specification (the SAML parties) and the end user using a web browser client (the client). Additionally, adversary can create and control any number of compromised agents. In fact, since honest agents only communicate over TLS, the adversary must create and use compromised agents to interact with the system as described in Section III-B.
@@ III-C System setup
System setup rules model the agent initialization, key generation, role assignment, resource registration, and trust establishment. The rules are general and for allow arbitrary complex system and trust relationships. A state machine illustrating the system setup rules is given in Fig. III-C. We assume unbounded number of agents (SPs, IdPs, Clients) as well as an unbounded of parallel sessions for each agent.
III-D SAML parties – the SP and the IdP
@@
As an example, in Listing 2 we provide the rule modelling the behavior of SP when receiving a resource request from an honest agent or adversary.
Client creates new TLS connection with unique tls_session_id and sends requests for resource located at url at $SP$. On the other side, SP awaits such request if other facts are already created:
  • !SP_Initialized(\$SP) — SP is initialized.
@@
  • Is_Resource(\$url, \$SP) — resource was registered during system setup (possibly by the adversary).
@@
  • !Is_Registered(\$SP, \$IdP) — IdP and SP have established trust during system setup.
@@
  • Fr(~rs) — fresh variable denoting relay state is created in this rule step.
@@
  • New_Id(~id_AuthnRequestuest) — precondition modelling creation of a fresh variable for to use as the message id of Authentication Request.
  • New_Server_TLS_Session(
    \$SP, ~tls_session_id_C_SP)
    - newly created TLS channel between C and SP.
Once SP receives the ResourceRequest and all the above facts are available, the SP is creates an authentication request <AuthnRequest>, and sends it along with RelayState back to client.
@@
As mentioned in section III-A, <AuthnRequest>, RelayState, and the rule creating the the term containing the message id — New_Id(~id_AuthnRequestuest) — vary according to the protocol variant. For example in the “strongid” variants the term is created by the rule that keeps the ~id_AuthnRequestuest secret, while and in “weakid” variants ~id_AuthnRequestuest is sent to network and, hence, made known to adversary.
1rule SP_Resource_Request:
2  let
3    // Received
4    ResourceRequest = <’ResourceRequest’, \$url>
5    Req = https_request(
6        \$SP, ~tls_session_id_C_SP, ResourceRequest)
7    // Building
8    AuthnRequestuest = <’AuthnRequest’,
9        ~id_AuthnRequestuest, \$SP, \$IdP>
10    OptionalSignature =
11        <#AUTHN_SIGNATURE AuthnRequestuest | ~ltkSP>
12    RelayState = <#RELAY_STATE_MESSAGE ~rs | \$url>
13    // Sending
14    Resp = https_response(~tls_session_id_C_SP,
15        <\$IdP, RelayState, AuthnRequestuest,
16            OptionalSignature>)
17  in
18  [
19    !SP_Initialized(\$SP),
20    !Is_Resource(\$url, \$SP),
21    !Is_Registered(\$SP, \$IdP),
22    !Ltk(\$SP, ~ltkSP),
23    New_Server_TLS_Session(\$SP, ~tls_session_id_C_SP),
24    In(Req),
25    New_Id(~id_AuthnRequestuest),
26    Fr(~rs)
27  ]
28  --[
29    Start_SP(\$SP),
30    SP_Resource_Request(\$SP, \$IdP, ~id_AuthnRequestuest,
31        \$url),
32    Https_Response(Resp)
33  ]->
34  [
35    SP_State_1(\$SP, \$IdP, AuthnRequestuest,
36        <#SP_SESSION_STATE ~rs | \$url>),
37    Out(Resp)
38  ]
Listing 2: Example Tamarin prover rule - SP recieves resource request
III-E Client modeling
@@
In our work, the focus has been on thoroughly modelling the system parts defined by the SAML technical standards — namely the SPs and IdPs. Since behaviour of the user and web client is not defined by SAML technical standards, we build the canonical user and the web client. We combine the end user and the web browser client into a single entity that we refer to as the client in our formal model and in this paper.
@@
The client is an idealized version of the the end user and it’s web browser client that does not take into account all the intricacies of the browser behaviour. For example, our client will not follow arbitrary redirects, will not store cookies etc. Also, in a more comprehensive model, one would need to take into account end user behaviour and possible human errors. For example, will the end user diligently verify the description of the service provider before authenticated and giving consent at the identity provider.
@@
During authentication to the IdP we assume that the end user will authenticate and consent only when the SP matches the service provider the client initiated the session with. Hence, we assume that the user is diligent enough to recognize the intended SP.
Moreover, the client will remember the SP it initialed the protocol with and only again connect to that specific SP when it received the artifact from the SP. In the real world, this is implemented using a combination of relay state and browser cookies.
Finally, client authentication to the IdP is also not part of the SAML standard, so we wrote our rules for client authentication. Before authentication occurs, client needs to be registered at IdP for witch we also wrote custom rules described in section III-C. Client has shared his $clientSecret$ (representing clients credentials) with IdP. While authenticating, after providing $clientSecret$ to the IdP over secure channel and after IdP successfully verifies the clients credentials, client is considered authenticated and IdP can now issue authentication assertion.
III-F The Adversary
As mentioned before, HTTPS requests and responses which are modeled by private functions and, hence, the adversary cannot construct them directly. Therefore, we introduce special rules that enable the adversary to send and receive HTTPS requests and responses, acting either as a server and as a client.
To participate in protocol flow, the adversary first needs to compromise some of honest agents and then he can communicate in their name with other honest agents. Two compromised agents are not able to communicate to each other. By implementing the adversary this way we enabled Tamarin built-in engine for the adversary which can create fresh facts, combine facts etc. Initialization rules that enable the adversary to pose for interaction with the protocol participants are marked with red arrows on Fig. III-C.
III-G Limitations
We make a number of simplifying assumptions to keep the complexity of the model manageable. We encountered some parts that are difficult to model in formal fashion or make the verification computationally unfeasible. Moreover, we needed to add some additional parameters for executability in formal model and for expressing security properties.
@@ III-G Limitations
Clients are a more idealized than SPs and IdPs as described in III-E.
@@
Rules for client communication with SP and IdP are basically describing the behaviour of a typical web browser. Aside from the initial request, other two client rules are just automatic web browser responses to HTTP scripts that are embedded as SPs and IdPs requests and responses. This could have been implemented in formal model by just passing the message from the input to the output, but because of formal model limitations, we implemented type checking of this automatic HTTP script requests and responses. Here we carefully designed rules so the client is not comparing values of message variables, just checking variable type (fresh, public etc.). Using type checking we removed significant number of meaningless message exchange through client automatic rules and therefore reduced Tamarin reasoning time. Of course, such type checking is not used in real code implementations and our model is, therefore, less faithful in that regard. Nevertheless, we argue that type checks are not a significant loss of generality, since the type checking is definitely happening at SP and IdP side. Moreover, for cases where the adversary would like to send something that does not correspond to message type, it can still do that through the adversary rules.
SAML SSO Profile known messages fields
SAML SSO Profile defines number of properties and message definitions, message ids, name-spaces, additional assertion parameters etc. These fields are well-specified in the protocol documentation and are not subject to modification or tampering by adversary. These properties are important in concrete implementations, but on formal level of abstraction they are irrelevant.
@@
Modeling too many execution paths can make the analysis intractable and the results difficult to interpret. Therefore we focused on the critical aspects and security-relevant behaviors of the POST/Artifact use case and did not model all possible execution paths in SAML SSO Profile that we modeled. Example execution paths that are not modeled are if client sends wrong request then SP returns error, if SP sends wrong format authentication request then IdP returns error, if client sends multiple resource request etc.
@@
For expressing security property of resource secrecy we needed to add additional parameters out of scope of SAML specification. While delivering resource to Client in final step, SP generates fresh unique variable which we called temporary secret $ts$. Without this session unique value we could not express property regarding the secrecy of resource delivered to client in specific session. We did the same with assertion issued by IdP to SP where we inserted assertion secret variable $sec$.
IV Security properties and analysis results
In this section, we describe the security properties we consider and discuss the results of the formal analysis.

### v2

@@ IV-A Methodology and Model Variants
¡AuthnRequest¿    ¡Subject¿ Specifies the subject of the authentication request or assertion, which identifies the principal that is the subject of the request or assertion.
¡AuthnRequest¿    ¡Version¿ Specifies the version of the SAML protocol.
¡AuthnRequest¿    ¡IssueInstant¿ Specifies time at which the request or assertion was issued.
¡AuthnRequest¿    ¡ForceAuthn¿ Indicates whether the identity provider must authenticate the principal even if the principal has an existing session with the identity provider.
¡AuthnRequest¿    ¡IsPassive¿ Indicates whether the identity provider must not interact with the principal, such as by prompting the principal for credentials, during the authentication process.
¡Response¿    ¡EncryptedAssertion¿ Represents an assertion that has been encrypted to protect its confidentiality.
¡Response¿    ¡Assertion¿     ¡AuthnStatement¿ Contains information about the authentication of a principal, including the time of authentication and the method of authentication used.
¡Response¿    ¡Assertion¿     ¡Subject¿      ¡SubjectConfirmationData¿       ¡OneTimeUse¿ Indicates whether the authentication request or assertion can only be used once, to prevent replay attacks.
¡Response¿    ¡Assertion¿     ¡Subject¿      ¡SubjectConfirmationData¿       ¡NotOnOrAfter¿ Specifies the time after which the authentication request or assertion is no longer valid.
@@
Table II: Elements that are not implemented in our model which are required by SAML standard or/and relevant to the security properties of protocol.
As mentioned before, there are several degrees of freedom in the SAML specifications, and here we describe the methodology by which we converted the specifications to the core formal model of SAML and well as the eight protocol variants we consider in our analysis.
Our approach for optional features was as follows:
  • All required fields and processing rules (i.e., those described with MUST key word in the standard) are included in the core model, with the following exception: If the lack of processing rules gives strong evidence that the field plays a purely functional rather than security role, we leave it out of the model.
Example required fields not included in the model are the <Version> and the <Timestamp> of the <Assertion>. An example sensitive optional feature is the use of <EncryptedAssertion> instead of <Assertion> in authentication responses. This feature is clearly security relevant, and we recognize as such. Note, however, we leave the implementation for future work rather than adding it to the core model right now. Table II contains the exhaustive list of all fields of <AuthnRequest> and <Response> not included in our core model, with the fields we consider security sensitive.
The other source of difficulty when converting the standard to the formal model is the lack of specification of various mechanism. Here, we tried to assess if the under-specified mechanism plays a purely functional role or if it could potentially have an impact on the security properties of the overall system. In the latter case, we include the mechanism in the list of sensitive under-specified mechanisms. For example, the SAML standard specifies that “dereferencing of the artifact using the Artifact Resolution profile MUST be mutually authenticated, integrity protected, and confidential“ [22], but does not explicitly specify the mechanisms to achieve these properties.
When deciding how exactly to model an under-specified mechanism, we considered a straightforward canonical implementation and, additionally, examined various SAML implementations to identify how are those under-specified mechanisms implemented in practice.
After identifying the sensitive optional features and sensitive under-specified mechanisms, we decided to pick three of these and model each in two different ways, yielding eight different protocol variants.
Even with automated formal analysis, it was challenging to model and verify all possible protocol variants (specific challenges are discussed in Section V-F). In this paper, we opted to focus on three specific degrees of freedom, described below. Since one of our goals was to investigate if, and how, small changes in the interpretation of the standard impact the security of the complete system, we picked the features and implementations that seemed most likely to have a significant impact, but are also illustrative of the kinds of choices and potential pitfalls in building a standards-compliant SAML deployment. We acknowledge that, for a truly comprehensive analysis, it is necessary to consider all properties from Table II, and even elements beyond the scope of SAML specification such as the user behavior. However, such extensions are reserved for future work.
@@
SAML specification mandates the use identifiers for many SAML messages. The specification requires that “any party that assigns an identifier MUST ensure that there is negligible probability any other party will accidentally assign the same identifier to a different data object[29]. Notice that this does not imply that identifiers must be generated via a cryptographic pseudorandom number generator (PRNG) — only that the probability of collisions between identifiers generated by honest protocol parties is negligibly low. Clearly, this is a sensitive under-specified mechanisms since SAML identifiers are used to match responses to the original requests.
@@
We model two different behaviors corresponding to cryptographically strong and weak random identifiers:
@@ IV-B Communication channels
Our model of TLS communication is similar to one in [36], except that we use ordinary network channels for communication and wrap TLS messages inside private function names. Using ordinary network channels significantly reduced precomputation time while wrapping TLS messages inside private functions makes it impossible for adversary to read the messages. Furthermore, we do not require channel IDs as in [36] since TLS sessions in SAML SSO are short-lived (only one request and response).
TLS session is established between two agents (the notion of agents is precisely described in the following Section IV-C). For every TLS request towards an agent acting as server S, a new session is created and labeled with unique tls_session_id. Since we model one-way authenticated TLS, any agent can act as client and initiate such session. However, only the agent S can access the data in the request, and create a response bound to the same tls_session_id. Such response can be accessed only by the agent that initiated the TLS session. Note that the agent’s name S doubles as its TLS identity, in practice this can be, for example, its common name or a pinned certificate.
Figure 2: Prerequisites Rules state machine. Initialized agent can become one of three honest parties or become compromised. Clients (and the adversary) can register resources at SPs (honest or compromised). Client (and the adversary) can register at IdPs (honest or compromised). Finally, any two SAML parties (SP and IdP) can establish trust.
IV-C System setup
In our model, the system consists of a number of agents. There are three kinds of honest agents – the SPs and IdPs whose behavior is defined by the SAML specification (the SAML parties) and the end user using a web browser client (the client). Additionally, adversary can create and control any number of compromised agents. In fact, since honest agents only communicate over TLS, the adversary must create and use compromised agents to interact with the system as described in Section IV-B.
@@ IV-C System setup
Listing 1: Example Tamarin prover meta-rule — SP recieves resource request — annotated with information on the source and purpose of particular elements of the rule.
System setup rules model the agent initialization, key generation, role assignment, resource registration, and trust establishment. The rules are general and for allow arbitrary complex system and trust relationships. A state machine illustrating the system setup rules is given in Fig. 2. We assume unbounded number of agents (SPs, IdPs, Clients) as well as unbounded number of parallel sessions for each agent.
IV-D SAML parties – the SP and the IdP
@@
As an example, in Listing 1 we provide the rule modelling the behavior of SP when receiving a resource request from an honest agent or adversary (corresponding to the actions of the SP after receiving the very first request in Fig. 1). Technically, this is a meta-rule as preprocessor directives (namely, #AUTHN_SIGNATURE, #RELAY_STATE_MESSAGE and #SP_SESSION_STATE) are use to craft the exact rule for a given protocol variant.
This is the initial step of protocol for SP, so this rule’s pre-conditions contain facts modelling SP’s initial state. These facts are persistent — they can be consumed many times, corresponding to multiple sessions of the same SP.
  • !SP_Initialized(\$SP) — persistent fact denoting that SP has been initialized.
@@
  • !Is_Resource(\$url, \$SP) — persistent fact denoting that resource was registered at SP during system setup (possibly by the adversary).
@@
  • !Is_Registered(\$SP, \$IdP) — persistent fact denoting that SP and a certain IdP have established trust during system setup.
  • !Ltk(\$SP,~ltkSP) — persistent fact containing SPs private key created during SP initialization.
The Client has created a new TLS connection with unique TLS session ID in order to send request for resource located at \$url at SP. The TLS connection is denoted with the fact New_Server_TLS_Session(…) that SP consumes at this step. The client’s request $Req$ is delivered to this rule step through Tamarin built-in fact In(). SP will consume the In(Req) and create fresh values needed to build the authentication request that will be sent as response to client’s request. The creation of these fresh values is also a part of rule pre-conditions:
@@
  • Fr(~rs) — fresh value denoting relay state is created in this rule step.
@@
  • New_Id(~id_AuthnRequest) — precondition modelling creation of a fresh value to use as the SAML identifier of the authentication request.
Finally, SP sends the authentication request along with the relaystate as a response $Res$ using Tamarin builtin fact Out(), and creates the fact SP_State_1(…) which denotes that SP is now in state where it received request for resource located at \$url from some client and that SP has sent an authentication request via the client to the chosen IdP.
@@
As mentioned in Section IV-A, <AuthnRequest>, RelayState, and the rule creating the term containing the SAML ID — New_Id(~id_AuthnRequest) — vary according to the protocol variant. For example, in the “strongid” variants the term is created by the rule that keeps the ~id_AuthnRequest secret, while in “weakid” variants it is sent to the network and, hence, made known to the adversary.
IV-E Client modeling
@@
In our work, the focus has been on thoroughly modelling the system parts defined by the SAML technical standards — namely the SPs and IdPs. Since behavior of the user and web client is not defined by SAML technical standards, we build the canonical user and the web client. We combine the end user and the web browser client into a single entity that we refer to as the client in our formal model and in this paper.
@@
The client is an idealized version of the end user and its web browser client that does not take into account all the intricacies of the browser behavior. For example, our client will not follow arbitrary redirects, will not store cookies etc. Also, in a more comprehensive model, one would need to take into account end user behavior and possible human errors. For example, during authentication to the IdP we assume that the end user will authenticate and consent only when the SP matches the service provider the client initiated the session with. Hence, we assume that the user is diligent enough to recognize the intended SP.
Moreover, the client will remember the SP it initialized the session with, and only again connect to that specific SP when it received the artifact from the SP. In the real world, this is implemented using a combination of relay state and browser cookies.
@@
Client authentication to the IdP is also not part of the SAML standard, so we created rules modelling simple client authentication. Firstly, client registration at IdP during system setup (see Section IV-C) sets up a shared secret $clientSecret$ between the client and IdP. While authenticating, after providing $clientSecret$ to the IdP over secure channel and after IdP successfully verifies it, the client is considered authenticated and IdP continues the protocol flow.
IV-F Threat Model and The Adversary
The threat model used in this paper closely matches the SAML threat model as identified by the SAML standard documents [23]. The only notable difference is that we are not considering denial-of-service attacks — we feel that this is a complementary problem and not well suited for symbolic protocol analysis.
Tamarin prover comes equipped with the generic Dolev-Yao [30] adversary that completely controls the network and is able to eavesdrop and, for example, launch man-in-the-middle or replay attacks. However, we model HTTPS requests and responses by private functions and, hence, the generic Tamarin adversary cannot construct them directly. Therefore, we introduce special rules that enable the adversary to send and receive HTTPS requests and responses, acting either as a server and as a client.
As in the SAML threat model, we assume that there could be a number of dishonest colluding participants in the protocol. We model these by allowing the adversary to compromise arbitrary agents. When an agent is compromised, the adversary gains access to its long term keys and secrets, and act as an HTTPS server using the compromised agent’s identity.
Furthermore, the adversary can establish trust relationships between compromised and other, honest, agents. Our model allows for arbitrary relationships. For example, an adversary can control a compromised client $C$ that has a resource registered at an honest SP $S$, and user accounts at an honest IdP $I_{1}$ and at a compromised IdP $I_{2}$. Meanwhile, the SP $S$ could be configured to trust both IdPs $I_{1}$ and $I_{2}$. Initialization rules that enable the adversary to compromise protocol participants and setup trust relationships are denoted with red arrows in Fig. 2.
IV-G Limitations
We make a number of simplifying assumptions to keep the complexity of the model manageable. Most notable limitations are as follows.
@@ IV-G Limitations
Clients are more idealized (compared to SPs and IdPs) and include both the end user and the web browser, as described in Section IV-E. This simplified client model limits the scope of our analysis — we will not discover attacks that arise from certain client-side behavior, such as browser-specific details (e.g., automatic redirects, cookie management), user errors (e.g., phishing or UI-based attacks), or the combination of these with the SAML protocol itself. Similarly, abstracting client authentication to IdP may impact our analysis by failing to discover attacks associated with the specific client authentication mechanisms.
@@
Rules for client communication with SP and IdP are describing the behavior of a typical web browser. Aside from the initial request, the other two client rules are just automatic web browser responses to HTTP scripts that are embedded as SPs and IdPs requests and responses. This could have been implemented in formal model by just passing the message from the input to the output. In our model, however, we implemented type checking of messages forwarded by clients. These rules are carefully designed, so the client is not comparing values of message variables, just checking variable type (fresh, public etc.). Using type checking, we removed a significant number of message exchanges through client automatic rules and, therefore, reduced Tamarin reasoning time. Of course, such type checking is not used in real code implementations and our model is, therefore, less faithful in that regard. Nevertheless, we argue that type checks are not a significant loss of generality, since these messages are again type checked at SP and IdP side. Moreover, for cases where the adversary would like to send something that does not correspond to the message type, it can still do that through the adversary rules.
SAML SSO Profile messages fields
SAML SSO Profile defines a number of properties and message definitions, message ids, name-spaces, additional assertion parameters etc., that we do not include in the formal model. These fields are well-specified in the protocol documentation and important in concrete implementations, but are not relevant to the security properties we consider in this paper.
@@
Modeling too many execution paths can make the analysis intractable and the results difficult to interpret. Therefore, we focused on the critical aspects and security-relevant behaviors of the POST/Artifact use case and did not model all possible execution paths. Example execution paths that are not modeled are related to various error conditions (e.g., client sends an incorrect request).
@@
In order to express all desired security properties (most notably the resource secrecy), we needed to add additional parameters not present SAML specification. While delivering resource to Client in the final step, SP generates a fresh unique variable which we called temporary secret $ts$. Without this session unique value, we could not express property regarding the secrecy of resource delivered to the client in specific session. We did the same with assertion issued by IdP to SP where we inserted assertion secret variable secsec.
V Security properties and analysis results
Now we present the security properties we consider in our analysis, their formalization as Tamarin lemmas, and the results of the automated verification.
V-A Formalization and verification of security properties
Listing 2 presents two Tamarin lemmas — one corresponding to the client resource authenticity property and one corresponding to the strong variant of client resource secrecy property — we include them as representative examples to illustrate how security goals are expressed and verified within our formal model. As mentioned in Section III-D, we use exclusively trace properties, which are expressed as guarded first-order logic formulas. The lemma is true if the formula holds for all possible traces (executions) of the protocol.
1lemma sec_Client_Resource_Authenticity:
2"All C S I cid csid url res ts #i.
3 Client_Params(C, S, I, cid, csid, <res, url, ts>)@i
4 ==>
5 (Ex #j. Register_Resource_at_SP(C, S, cid, res, url)@j)
6 | (Ex #j. Compromise_Agent(S)@j)
7 | (Ex I1 #j #k. Compromise_Agent(I1)@j
8 & Register_SP_at_IdP(S, I1)@k)
9 "
10
11lemma sec_Client_Registered_Resource_Secrecy_Strong:
12 "All C S url res id #i.
13 Register_Resource_at_SP(C, S, id, res, url)@i
14 ==>
15 (not (Ex #j. K(res)@j))
16 | (Ex #j. Compromise_Agent(S)@j)
17 "
Listing 2: Example lemmas
In both lemmas, the variables i, j, k are typed as timestamps, where other variables are untyped. Predicates Client_Params, Register_Resource_at_SP, etc. correspond to the action facts of the rules modeling SAML participants, while the Compromise_Agent corresponds to the setup rule that allows the adversary to compromise agents. The predicate K denotes the knowledge of the adversary.
Informally, the client resource authenticity lemma states the following: if a client C (at some timestamp i) completes a protocol session with the SP S and the IdP I with session parameters client ID cid, obtaining resource res at URL url (line 3), then C has (at some timestamp j) registered that exact resource res located at url at the SP S (line 5). Furthermore, the lemma allows that C has not registered the resource at the SP S if one of the following holds: the SP S is compromised (line 6), or there is some compromised IdP I1 that SP S trusts (line 7).
The strong client registered resource secrecy lemma states the following: if a client C registers a resource res at the SP S (line 13) then that resource remains unknown to the adversary — it is never the case that the adversary knows the resource res at some timestamp j (line 15). Furthermore, the lemma allows that adversary learns res if the specific SP S is compromised (line 16).
Hence, in this lemma, we ask that secrecy holds even if there exists some compromised IdP I’ trusted by SP. This lemma is included for sanity checking purposes as it should clearly fail — a compromised IdP I’ trusted by S can initiate a session, craft authentication responses for any client of its choice, and obtain their resources. Formal analysis validates this intuition and confirms that this lemma is falsified by an expected attack trace.
Note that the lemmas allow us to reason not only about the actions of participants in the protocol, and the knowledge of the adversary, but also express fine-grained honesty assumptions on the protocol participants by reasoning about events happened during the trust setup.
Tamarin proves trace properties by attempting to falsify them — given a lemma, it tries to construct a trace that satisfies the negation of the property. If such a counterexample is found, it represents a concrete attack. If no such trace exists and the search terminates, Tamarin produces a formal proof in the form of a proof tree, demonstrating that the property holds for all possible traces. Since verifying trace properties is undecidable in general, Tamarin may not always terminate. In our analysis, we encountered several lemmas for which automated verification did not terminate or was prohibitively slow. To address this, we restructured the model to reduce ambiguity, and introduced helper lemmas to capture useful invariants, most notably related to TLS session and long term secrets. Additionally, we created proof-guiding oracles to direct the prover by prioritizing proof goals. These were crafted based on the insights gained from attempts at manual verification of lemmas using the Tamarin’s interactive mode.
V-B Overview of security properties and analysis results
On a high level of abstraction, a Web SSO protocol should protect against unauthorized access and ensure confidentiality of sensitive user data. The exact desired security properties that the SAML Web SSO Profile aims to achieve are not explicit or precise in the official SAML documentation. The “Security considerations” non normative document [23] contains high-level descriptions of the threat model and the list of security techniques used to achieve specific low-level goals. For the SAML Web SSO profile, the standard [23] identifies several specific threats and possible countermeasures, both within and outside the standard. For example, one identified threat is the “Theft of the Bearer Token“ — the adversary that obtains an artifact will be able to impersonate the user. The countermeasures for this threat include using TLS to secure the connection, which is mandated by the standard. Finally, [23] gives multiple practical guidelines for implementing systems based on the SAML framework in the secure manner.
Compared to SAML documentation, we consider a broader set of security properties and offer their precise formalization based on usual patterns used for formally modeling and verifying security protocols [46, 34, 35, 25]. The properties considered in this paper encompass most of the threats described in [23] with denial-of-service attacks being the notable exception.
Most properties we consider can be grouped into three categories:
  • Secrecy properties state that the adversary cannot obtain information that is supposed to stay secret (e.g., the end-user resource or the details of the assertion). In terms of [23], these correspond to threats such as exposure of artifacts or assertions. Note that our secrecy properties are much more general than just the robustness against eavesdropping. For example, the adversary can attempt to compromise resource secrecy by launching man-in-the middle attacks, compromising various agents etc.
  • Agreement (or authentication) properties state that the views of honest agents participating in the protocol session are compatible to some degree. Agreement properties differ when it comes to the details of what exactly has to match in the compatible views of two protocol agents, and we consider all variants of authentication properties described in the seminal paper [46] — namely weak, non-injective and injective agreement. In terms of [23], these properties imply that the forged assertions, message insertion, message deletion, message modification, or man-in-the middle attacks cannot happen in the protocol.
  • Freshness properties state that certain actions (such as user authentication) must have happened recently in the past. In terms of [23], these properties cover the resistance to replay attacks.
Finally, as shown in table, properties can be grouped by the participant from whose point of view we consider the protocol execution (client, SP and IdP) and by the exact assumptions on entities compromised by the adversary.
In total, we have considered 28 security properties and have either formally verified or falsified every property for each of the 8 protocol variants. The complete overview of the security properties and verification results is given in Table III. The rest of this section gives an overview of the most relevant properties.